Consider all samples of size of 5 from the population

2,5,7,9,10,11,12


1
Expert's answer
2021-07-06T14:37:24-0400

N=7

n=5

Number of samples "= \frac{N!}{n!(N-n)!}"

"= \frac{7!}{5!(7-5)!}= \frac{6 \times 7}{2}= 21"